A sum of money is shared among nine people
so that the first gets 75 naira next 150 naira
the next 225 naira,and so on
A.How much money does the ninth person get
B.How much money is shared altogether

Found 2 solutions by mananth, ikleyn:
Answer by mananth(16946)


A sum of money is shared among nine people
so that the first gets 75 naira next 150 naira
the next 225 naira,and so on
A.How much money does the ninth person get
B.How much money is shared altogether
The money shared is an arithmetic sequence
tn = a+(n-1)d
a= 75, n=9 , d=75
The 9th person will get
t9 = 75+(9-1)*75
tn = 75+600
tn = 675

ninth person gets 675 naira
Sn sum shared is thr total amount shared

Sn = n/2(a+l)
n=9, a =75, l= 675
sn = 9/2 (75+675)
sn = 9/2 * 750
Sn = 9*375 = 3375
Total amount 3375 naira

Answer by ikleyn(52787)

.

(A)  The 9-th person gets  75*9 = 675 naira.


(B)  Altogether  9 persons get  75*(1+2+3+4+5+6+7+8+9) = 75*45 = 3375 naira.

Solved.